OpenStudy (anonymous):

I don't want the answer given to me, but I'd like help finding it because math makes no sense to me at all..

OpenStudy (anonymous):



OpenStudy (phi):

can you factor the top and bottom ?

OpenStudy (anonymous):

I learned how with smaller equations, but the longer ones, I don't understand what to do.

OpenStudy (phi):

to factor x^2 -x -20 you do the following: 1) look at the last number: -20 the minus means the two "roots" will have different signs the minus sign on -x , short for -1 x, means the larger root is negative. 2) list all pairs of numbers that multiply to give you 20 1,20 2,10 4,5 using rule 1) make the larger number - and the smaller number + 1-20= -19 2-10= -8 4-5= -1 3) pick out the pair that matches -1 x (the middle term of x^2-x-20) the pair you want is 4 and -5 4) the factors are (x+4)(x-5)

OpenStudy (phi):

to factor x^2+8x+16 1) the + sign on +16 means both roots are the same sign the + sign on +8x means the larger root is +. because the roots have the same sign, the smaller root is also + 2) list the pairs of numbers that multiply to give you 16 1,16 2,8 4,4 3) using the rule in (1), both roots are + 1+16= 17 2+8= 10 4+4= 8 4) pick the pair that gives +8 to match +8x the pair you want is 4 and 4 5) the factors is (x+4)(x+4)

OpenStudy (phi):

you get \[ \frac{(x+4)(x-5)}{(x+4)(x+4)} = \frac{(x-5)}{(x+4)} \text{ x≠-4 }\]